OpenStudy (akgorgeous):

URGENT ALGEBRA QUESTION- For the function f(x) = x + 4, what is the ordered pair for the point on the graph when x = 3p?

OpenStudy (danjs):

The function is y = f(x) = x + 4 when x = 3p the y value will be y = f(3p) = 3p + 4 the ordered pair (x , y) is then (3p , 3p +4)